Erie Company manufactures a mobile fitness device called the Jogging Mate. The company uses standards to control its costs. T(For requirements 3 through 5, indicate the effect of each variance by selecting F for favorable, U for unfavorable, and
Answer #1
1 Standard labor hours allowed 8820 =19600*(27/60)
2 Standard labor cost allowed 51156 =8820*5.80
3 Labor spending variance 1988 U =53144-51156
4 Labor rate variance 1898 F =53144-(9490*5.8)
Labor efficiency variance 3886 U =5.8*(9490-8820)
5 Variable overhead rate variance 4745 U =45552-(9490*4.3)
Variable overhead efficiency variance 2881 U =4.3*(9490-8820)